OpenStudy (anonymous):

1. Is the equation true, false, or open? Explain. 9p + 8 = 10p + 7 (1 point) Open, there is a variable. True, the expressions are the same for all values of the variables. False, the expressions are never the same. 2. Is x = 1 a solution of the equation 2 – 8x = –6? (1 point) yes no

OpenStudy (cwrw238):

1. Have you tried to solve this equation?

OpenStudy (cwrw238):

subtract 9p form both sides: 9p - 9p + 8 = 10p - 9p + 7 8 = p + 7 subtract 7 from both sides 8-7 = p + 7 - 7 1 = p so p = 1 is the solution